# MMR-Mamba: Multi-Contrast MRI Reconstruction with Mamba and Spatial-Frequency Information Fusion

### v1

@@ Abstract
Multi-contrast MRI acceleration has become prevalent in MR imaging, enabling the reconstruction of high-quality MR images from under-sampled k-space data of the target modality, using guidance from a fully-sampled auxiliary modality. The main crux lies in efficiently and comprehensively integrating complementary information from the auxiliary modality. Existing methods either suffer from quadratic computational complexity or fail to capture long-range correlated features comprehensively. In this work, we propose MMR-Mamba, a novel framework that achieves comprehensive integration of multi-contrast features through Mamba and spatial-frequency information fusion. Firstly, we design the Target modality-guided Cross Mamba (TCM) module in the spatial domain, which maximally restores the target modality information by selectively absorbing useful information from the auxiliary modality. Secondly, leveraging global properties of the Fourier domain, we introduce the Selective Frequency Fusion (SFF) module to efficiently integrate global information in the frequency domain and recover high-frequency signals for the reconstruction of structure details. Additionally, we present the Adaptive Spatial-Frequency Fusion (ASFF) module, which enhances fused features by supplementing less informative features from one domain with corresponding features from the other domain. These innovative strategies ensure efficient feature fusion across spatial and frequency domains, avoiding the introduction of redundant information and facilitating the reconstruction of high-quality target images. Extensive experiments on the BraTS and fastMRI knee datasets demonstrate the superiority of the proposed MMR-Mamba over state-of-the-art MRI reconstruction methods.
@@
MRI reconstruction, Multi-contrast, State space models, Fourier domain, Spatial-frequency information fusion.
@@ I Introduction
Magnetic resonance imaging (MRI) is an essential clinical imaging technology owing to its non-invasive, free-of-radiation characteristics and its capability to provide high-resolution morphological information with varying contrast [1]. However, due to the intrinsic physics of MR imaging systems [2], acquiring fully sampled k-space data can take tens of minutes, resulting in patient discomfort and increased operational costs. Therefore, accelerating MRI acquisition, particularly through the reconstruction of high-quality MR images from undersampled k-space data has become a popular research topic [3].
In clinical practice, multi-contrast MR images are simultaneously acquired to enable more accurate disease diagnosis and treatment planning. These modalities provide complementary information, revealing detailed anatomical structures of the patient [4]. For instance, T1 and T2 weighted images (T1WIs and T2WIs), along with proton density and fat-suppressed proton density-weighted images (PDWIs and FS-PDWIs), are two pairs of images with complementary structures. Typically, T1WIs are acquired more quickly due to shorter repetition time (TR) and echo time (TE) requirements, whereas T2WIs acquisition requires longer TR times, resulting in slower imaging speeds [5]. Similarly, FS-PDWIs necessitate longer scan times than PDWIs. In this context, leveraging readily obtainable modalities (i.e. auxiliary modalities such as T1WIs or PDWIs) as supplementary guidance for the reconstruction of target modalities (e.g., T2WIs or FS-PDWIs) with slower imaging speeds, known as multi-contrast MRI reconstruction, has been investigated and verified by previous works. For instance, compressed sensing (CS), Bayesian learning, dictionary learning, and graph representation theory have been utilized to accelerate multi-contrast MR imaging [6, 7, 8]. More recently, deep learning has significantly advanced MRI reconstruction due to their powerful feature representation capabilities [5, 9, 7, 10, 11].
@@
In multi-contrast MR images, similar structural features are distributed across different regions within each modality, while complementary information is provided by different modalities. Consequently, the primary challenge of the multi-contrast MRI reconstruction task is to comprehensively explore long-range dependencies within each modality and effectively leverage complementary information from auxiliary modalities. CNN-based methods have been proposed to integrate multi-contrast information [5, 12], but they typically demonstrate local sensitivity and a lack of long-range dependency, thereby limiting their ability to capture global contextual information. In contrast, Transformer-based models [4, 13, 14], distinguished by their large receptive fields and global sensitivity, often surpass CNNs in capturing extensive contextual information. However, these models are burdened by substantial computational overhead due to the quadratic growth of resources with respect to sequence length. Therefore, developing an algorithm that comprehensively integrates complementary information from different modalities without significant computational overhead is crucial.
@@
Recently, the improved structured state-space sequence model with a selective scanning mechanism, Mamba [15], has emerged as a compelling alternative to Transformer, due to its ability to model long-range sequence relationships with linear complexity. Mamba has shown superior performance compared to Transformers in tasks involving long-term dependency modeling, such as natural language processing [16, 17] and medical image segmentation [18, 19]. Investigating Mamba’s potential for long-range dependency modeling and complementary information fusion in multi-contrast MRI reconstruction is highly promising. On the other hand, each component in the frequency domain represents a combination of all the pixel values in the spatial domain, meaning that frequency features capture the overall patterns and structures, providing a global view of the entire image. Meanwhile, Fourier features are beneficial for recovering high-frequency signals that are crucial for addressing image degradation[20]. Therefore, comprehensive and efficient global feature integration across different modalities can be achieved by performing feature fusion in the frequency domain. The Mamba block and frequency domain thus offer two promising solutions for efficient fusion of the multi-contrast information.
@@
Motivated by the above analysis, we propose a novel framework, MMR-Mamba, for multi-contrast MRI reconstruction. Built upon the Mamba architecture, our MMR-Mamba jointly explores the complementary information fusion in the spatial and frequency domains, implemented by the Target modality-guided Cross Mamba (TCM) module and Selective Frequency Fusion (SFF) module, respectively. Additionally, we introduce an Adaptive Saptial-Frequency Fusion (ASFF) module to enhance and integrate the fused features from both domains. Specifically, we first adapt Mamba blocks for feature extraction, leveraging their capability to effectively capture fine-grained relationships and long-range dependencies within MR images. Then we design TCM for spatial domain information fusion, where correlated features from the auxiliary modality are selectively supplemented to the target modality. In the SFF module of the frequency domain, we perform element-wise summation for the phase spectrum and selective integration for the amplitude spectrum, as the phase spectrum of both modalities primarily contains consistent structural information while the amplitude spectrum from different modalities holds incompatible style information. Finally, we adopt the ASFF module to enhance the fused features from both domains, where less informative channels from one domain are supplemented by incorporating the corresponding channel features from the other domain. The ASFF module enables the integration of relevant information and the suppression of redundant features. Our contributions can be summarised as follows:
@@
  • We propose MMR-Mamba, an efficient framework for multi-contrast MRI reconstruction. To the best of our knowledge, this is the first exploration of Mamba for integrating complementary information across multi-contrast MR images.
@@
  • We introduce the ASFF module for spatial-frequency information fusion, enhancing task-relevant features while suppressing irrelevant features from the two domains.
  • Extensive experiments on the BraTS and fastMRI datasets validate the effectiveness of our MMR-Mamba framework, demonstrating superior performance compared to existing methods.
@@ II-A Spatial Domain MRI Reconstruction
Benefiting from the complementary information from multi-contrast images, many methods have been proposed to reconstruct high-quality images from low-quality images in the spatial domain, guided by the auxiliary modality images[21, 5, 4, 22]. Previously, Xiang et al.[5] proposed Dense-Unet to reconstruct the T2WIs from the T1WIs and under-sampled T2WIs, with the concatenation of under-sampled T2WIs and T1WIs serving as the input of the network. Xuan et al. [12] further enhanced the performance by introducing a spatial alignment network to compensate for spatial misalignment between multiple modality images.
Recently, regarding the Transformer’s ability in modeling long-range dependencies, several works have explored Transformer-based approaches for multi-contrast MRI reconstruction. MTrans [4] employed a cross-attention module for multi-scale features fusion of the target modality and the auxiliary modality. MD-GraphFormer [22] incorporated the physical constraint of MRI into the network architecture, and modeled the multiple contrasts as graph nodes for joint reconstruction of multi-contrast MR images over the graph. MCCA [13] introduced the hybrid convolutional transformer operation to enrich global and local spatial context representation. DCAMSR [14] proposed a dual cross-attention mechanism, where the features of the reference image and the upsampled input image are extracted and promoted with both spatial and channel attention at multiple resolutions. Despite the promising performance, the CNN and Transformer-based approaches either fall short of fully capturing long-range dependencies or suffer from quadratic complexity. In contrast, our model achieves comprehensive exploration of long-range dependencies without imposing a heavy computational burden.
@@ II-B Dual-Domain MRI Reconstruction
However, current dual-domain methods typically use similar networks for both k-space and image domains, which overlooks the specific characteristics inherent to each domain. Moreover, these methods often rely on straightforward parallel or sequential combinations of two domains, failing to fully exploit their unique features and complementary information.
@@ II-C State Space Models
State Space Models (SSMs) offer a powerful framework for efficiently modeling long sequences and have garnered significant attention [31, 32]. Previously, Structured State Space Sequence Models (S4) [33] was designed to capture long-range dependencies within sequences by introducing Higher-Order Polynomial Project Operator [34]. S4 has demonstrated exceptional performance across various benchmarks while reducing computational complexity to $O(Nlog(N))$, significantly superior to the quadratic computational complexity $O(N^{2})$ of Transformers. More recently, S6, i.e. Mamba [15], was introduced. It employs a selection mechanism for choosing relevant information based on input and features an efficient hardware-aware algorithm grounded in selective scanning.
Mamba has significantly advanced natural language tasks, surpassing traditional Transformers. Subsequently, Mamba has been extended to various vision tasks. For instance, Visual Mamba (Vim) [35] proposed a pure Mamba-based backbone utilizing bidirectional Mamba blocks, modeling images in a sequence manner through positional embedding, akin to ViT [36]. Visual State Space Model (Vmamba) [37] introduces a cross-scanning mechanism for spatial traversal, converting non-causal visual images into ordered patch sequences. In the medical imaging domain, Mamba has been explored for tasks such as medical image segmentation [18, 19], registration [38], and classification [39, 40]. For MRI reconstruction, Huang et al.[41] first proposed MambaMIR with an Arbitrary-Mask mechanism for medical image reconstruction. However, the complementary information from muti-contrast images is neglected. In this work, we explore Mamba for the integration of features from different modalities, harnessing its capability in modeling long-range dependencies.
@@ III-A Preliminaries
Fourier Transform. We here briefly review the meaning of the Fourier transformation of images for a better understanding of our work. The Fourier transform (FT) serves as a crucial technique in analyzing the frequency characteristics of an image. Transforming images from the spatial domain to the frequency domain through FT allows us to examine the images from a global perspective. Given an image $x$, the FT can be expressed as follows:
@@
where $u$ and $v$ are coordinates in the Fourier space. The frequency domain feature $\mathcal{F}(x)$ is represented as $\mathcal{F}(x)=\mathcal{R}(x)+j\mathcal{I}(x)$, with $\mathcal{R}(x)$ and $\mathcal{I}(x)$ denote the real and imaginary part respectively. Then the amplitude spectrum $\mathcal{A}(x)(u,v)$ and phase spectrum $\mathcal{P}(x)(u,v)$ are defined as:
@@
According to the Fourier theory, the amplitude spectrum $\mathcal{A}$ encodes the style information of an image, whereas the phase component spectrum $\mathcal{P}$ characterizes the structure information [42, 43]. Furthermore, as highlighted in [44, 45] the phase spectrum maintains global structure well and the degradation of images is mainly manifested in the amplitude spectrum.
@@ III-B Overall Architecture
Let ${I}_{tar}$ and ${I}_{aux}\in\mathbb{R}^{H\times W}$ be a pair of MR images from the target modality and auxiliary modality. Our objective is to reconstruct the high-quality target modality image $I_{tar}$ through a network that takes the under-sampled target modality image $\widetilde{I}_{tar}$ and fully-sampled auxiliary modality image ${I}_{aux}$ as inputs. Here, $\widetilde{I}_{tar}=\mathcal{F}^{-1}(\mathcal{M}\odot k_{tar})$, where $\mathcal{F}^{-1}$ represents the inverse Fourier transform (IFT), $\mathcal{M}$ denotes the binary under-sample mask, and $k_{tar}$ represents the fully-sampled k-space data of the target modality. The multi-contrast reconstruction task can be formulated as the following optimization problem:
$\hat{I}_{tar}=\operatorname*{argmin}_{f(\theta)}\|I_{tar}-f(\mathcal{F}^{-1}(\mathcal{M}\odot k_{tar}),I_{aux};\theta)\|_{1}.$ (6)
To fully harness and fuse the information from different modalities, we design the MMR-Mamba network, depicted in Fig. 1. Initially, we adopt CNN blocks $\psi$ for low-level feature extraction and Mamba blocks $\phi$ for high-level feature extraction, producing modality-specific features $F_{tar}$ and $F_{aux}$ for the target modality and auxiliary modality, respectively.
$F_{tar}=\phi_{tar}(\psi_{tar}(\widetilde{I}_{tar})),~{}F_{aux}=\phi_{aux}(\psi_{aux}({I}_{aux})).$ (7)
Subsequently, we design the Target modality-guided Cross Mamba (TCM) module in the spatial domain and Selective Frequency Fusion (SFF) module in the frequency domain, generating domain-specific fused features $F_{spa}$ and $F_{fre}$:
$F_{spa}=TCM(F_{tar},F_{aux}),~{}F_{fre}=SFF(F_{tar},F_{aux}).$ (8)
Lastly, the multi-contrast features in these two domains are further integrated through an Adaptative Spatial-Frequency Fusion (ASFF) module to generate the final fused feature, which is then fed into a CNN Decoder to reconstruct the target modality image:
@@
The model is trained by minimizing the $L_{1}$ loss between the reconstructed image and ground-truth image:
$L=\frac{1}{N}\sum_{n=1}^{N}\|\hat{I}_{tar}-I_{tar}\|_{1},$ (10)

### v2

# MMR-Mamba: Multi-Modal MRI Reconstruction with Mamba and Spatial-Frequency Information Fusion

@@ Abstract
Multi-modal MRI offers valuable complementary information for diagnosis and treatment; however, its utility is limited by prolonged scanning times. To accelerate the acquisition process, a practical approach is to reconstruct images of the target modality, which requires longer scanning times, from under-sampled k-space data using the fully-sampled reference modality with shorter scanning times as guidance. The primary challenge of this task is comprehensively and efficiently integrating complementary information from different modalities to achieve high-quality reconstruction. Existing methods struggle with this: 1) convolution-based models fail to capture long-range dependencies; 2) transformer-based models, while excelling in global feature modeling, struggle with quadratic computational complexity. To address this dilemma, we propose MMR-Mamba, a novel framework that thoroughly and efficiently integrates multi-modal features for MRI reconstruction, leveraging Mamba’s capability to capture long-range dependencies with linear computational complexity while exploiting global properties of the Fourier domain. Specifically, we first design a Target modality-guided Cross Mamba (TCM) module in the spatial domain, which maximally restores the target modality information by selectively incorporating relevant information from the reference modality. Then, we introduce a Selective Frequency Fusion (SFF) module to efficiently integrate global information in the Fourier domain and recover high-frequency signals for the reconstruction of structural details. Furthermore, we devise an Adaptive Spatial-Frequency Fusion (ASFF) module, which mutually enhances the spatial and frequency domains by supplementing less informative channels from one domain with corresponding channels from the other. Extensive experiments on the BraTS and fastMRI knee datasets demonstrate the superiority of our MMR-Mamba over state-of-the-art reconstruction methods.
@@
MRI reconstruction, multi-modal, State space models, Fourier domain, Spatial-frequency information fusion.
@@ I Introduction
Magnetic resonance imaging (MRI) is an essential clinical imaging technology due to its non-invasive, free-of-radiation characteristics and its capability to provide high-resolution morphological information with varying modalities [1]. In clinical practice, multi-modal MR images with complementary information are simultaneously acquired to enable more accurate disease diagnosis and treatment planning[2]. For example, in brain MR imaging, T1 weighted images (T1WIs) provide detailed anatomical structure information, while T2 weighted images (T2WIs) are useful for detecting edema, inflammation, and fluid-filled structures [3]. Similarly, in knee imaging, proton density weighted images (PDWIs) reveal structural information while fat-suppressed proton density weighted images (FS-PDWIs) can suppress fat signals and highlight cartilage ligaments [4]. However, MR imaging is inherently time-consuming due to sequential data acquisition in k-space, which results in patient discomfort and increased operational costs [5]. Therefore, accelerating MR imaging, particularly through the reconstruction of high-quality MR images from under-sampled k-space data, is highly demanded in clinical practice [6].
@@
Previous studies [7, 8, 9, 10, 11, 12, 13] have demonstrated that it is a promising way to leverage readily obtainable modalities (i.e. reference modalities such as T1WIs or PDWIs) as supplementary guidance for the reconstruction of target modalities (e.g., T2WIs or FS-PDWIs) with slower imaging speeds, known as multi-modal MRI reconstruction. The primary challenge of this reconstruction task is to comprehensively explore long-range dependencies within each modality and effectively leverage complementary information from reference modalities. Early methods employ convolutional neural networks (CNNs) to integrate multi-modal information [10, 14, 15], but they typically demonstrate local sensitivity and a lack of long-range dependency, thereby limiting their ability to integrate correlated features from both modalities for faithful MRI reconstruction. In contrast, Transformer-based models [2, 16, 17], distinguished by their large receptive fields and global sensitivity, often surpass CNNs in capturing extensive contextual information. However, these models are burdened by substantial computational overhead due to the quadratic growth of resources with respect to sequence length. Therefore, developing an algorithm that can comprehensively explore long-range dependencies and integrate complementary information from different modalities without significant computational overhead is crucial.
@@
Recently, the improved structured state-space sequence model with a selective scanning mechanism, Mamba [18], has emerged as a compelling alternative to Transformer, due to its ability to model long-range sequence relationships with linear complexity. Mamba has shown superior performance compared to Transformers in tasks involving long-term dependency modeling, such as natural language processing [19, 20] and medical image segmentation [21, 22]. Investigating Mamba’s potential for long-range dependency modeling and complementary information fusion in multi-modal MRI reconstruction is highly promising. On the other hand, each component in the frequency domain represents a combination of all the pixel values in the spatial domain, meaning that frequency features capture the overall patterns and structures, providing a global view of the entire image. Meanwhile, existing studies indicate that Fourier features are beneficial for recovering high-frequency signals that are crucial for addressing image degradation[23]. Therefore, comprehensive and efficient global feature integration across different modalities can be achieved by performing feature fusion in the frequency domain. Thus, Mamba block and frequency domain offer two promising solutions for comprehensive and efficient fusion of multi-modal information.
@@
Motivated by the above analysis, we propose a novel framework, MMR-Mamba, for multi-modal MRI reconstruction. Built upon the Mamba architecture, our MMR-Mamba jointly explores complementary information fusion in the spatial and frequency domains, implemented by the Target modality-guided Cross Mamba (TCM) module and Selective Frequency Fusion (SFF) module, respectively. Additionally, we introduce an Adaptive Spatial-Frequency Fusion (ASFF) module to mutually enhance the features from these two domains. Specifically, Mamba blocks are employed to extract features from each modality. Then we design TCM for spatial domain information fusion, where correlated features from the reference modality are selectively supplemented to the target modality through Cross Mamba. In the SFF module of the frequency domain, we perform element-wise summation for the phase spectrum and selective integration for the amplitude spectrum, as the phase spectrum of both modalities primarily contains consistent structural information while the amplitude spectrum from different modalities holds incompatible style information. Finally, we adopt the ASFF module to enhance the fused features from both domains, where less informative channels from one domain are supplemented by incorporating the corresponding channel features from the other domain. The ASFF module enables the integration of relevant information and the suppression of redundant features. Our contributions can be summarised as follows:
@@
  • We propose MMR-Mamba, an efficient framework for multi-modal MRI reconstruction. To our knowledge, this is the first exploration of Mamba to integrate complementary information across multi-modal MR images.
@@
  • We introduce the ASFF module for spatial-frequency information fusion, enhancing task-relevant features while suppressing irrelevant features from these two domains.
  • Extensive experiments on the BraTS and fastMRI datasets validate the effectiveness of our MMR-Mamba framework, demonstrating its superior performance compared to existing methods.
@@ II-A Spatial Domain MRI Reconstruction
Benefiting from the complementary information from multi-modal images, many methods have been proposed to reconstruct high-quality images from low-quality images in the spatial domain, guided by the reference modality images[24, 10, 2, 25]. Previously, Xiang et al.[10] proposed Dense-Unet to reconstruct T2WIs from T1WIs and under-sampled T2WIs, with the concatenation of under-sampled T2WIs and T1WIs serving as the input of the network. Xuan et al. [15] further enhanced performance by introducing a spatial alignment network to compensate for spatial misalignment between multiple modality images.
Recently, due to the Transformer’s capability in modeling long-range dependencies, several works have explored Transformer-based approaches for multi-modal MRI reconstruction. MTrans [2] employed a cross-attention module for multi-scale feature fusion of the target modality and the reference modality. MD-GraphFormer [25] incorporated the physical constraint of MRI into the network architecture and modeled the multiple contrasts as graph nodes for the joint reconstruction of multi-modal MR images over the graph. MCCA [16] introduced the hybrid convolutional transformer operation to enrich global and local spatial context representation. DCAMSR [17] proposed a dual cross-attention mechanism, where the features of the reference image and the upsampled input image are extracted and promoted with both spatial and channel attention at multiple resolutions. Despite the promising performance, the CNN and Transformer-based approaches either fall short of fully capturing long-range dependencies or suffer from quadratic complexity. In contrast, our model achieves comprehensive exploration of long-range dependencies without imposing a heavy computational burden.
@@ II-B Dual-Domain MRI Reconstruction
However, current dual-domain methods typically use similar networks for both k-space and image domains, overlooking the specific characteristics inherent to each domain. Moreover, these methods often rely on straightforward parallel or sequential combinations of two domains, failing to fully exploit cross-domain complementary information.
@@ II-C State Space Models
State Space Models (SSMs) offer a powerful framework for efficiently modeling long sequences and have garnered significant attention [34, 35]. Previously, Structured State Space Sequence Models (S4) [36] was designed to capture long-range dependencies within sequences by introducing Higher-Order Polynomial Project Operator [37]. S4 has demonstrated exceptional performance across various benchmarks while reducing computational complexity to $O(Nlog(N))$, significantly superior to the quadratic computational complexity $O(N^{2})$ of Transformers. More recently, S6, i.e., Mamba [18], was introduced. It employs a selection mechanism for choosing relevant information based on input and features an efficient hardware-aware algorithm grounded in selective scanning.
Mamba has significantly advanced natural language tasks, surpassing traditional Transformers. Subsequently, Mamba has been extended to various vision tasks. For instance, Visual Mamba (Vim) [38] proposed a pure Mamba-based backbone utilizing bidirectional Mamba blocks, modeling images in a sequence manner through positional embedding, akin to ViT [39]. Visual State Space Model (Vmamba) [40] introduces a cross-scanning mechanism for spatial traversal, converting non-causal visual images into ordered patch sequences. In the medical imaging domain, Mamba has been explored for tasks such as medical image segmentation [21, 22], registration [41], and classification [42, 43]. For MRI reconstruction, Huang et al.[44] first proposed MambaMIR with an Arbitrary-Mask mechanism for medical image reconstruction. However, the complementary information from multi-modal images is neglected. In this work, we explore Mamba for the integration of features from different modalities, harnessing its capability in efficiently modeling long-range dependencies.
@@ III-A Preliminaries
Fourier Transform. Here, we briefly review the meaning of the Fourier transformation of images for a better understanding of our work. The Fourier transform (FT) serves as a crucial technique in analyzing the frequency characteristics of an image. Transforming images from the spatial domain to the frequency domain through FT allows us to examine the images from a global perspective. Given an image $x$, the FT can be expressed as follows:
@@
where $u$ and $v$ are coordinates in the Fourier space. The frequency domain feature $\mathcal{F}(x)$ is represented as $\mathcal{F}(x)=\mathcal{R}(x)+j\mathcal{I}(x)$, with $\mathcal{R}(x)$ and $\mathcal{I}(x)$ denote the real and imaginary part respectively. Then, the amplitude spectrum $\mathcal{A}(x)(u,v)$ and phase spectrum $\mathcal{P}(x)(u,v)$ are defined as:
@@
According to the Fourier theory, the amplitude spectrum $\mathcal{A}$ encodes the style information of an image, whereas the phase component spectrum $\mathcal{P}$ characterizes the structure information [45, 46]. Furthermore, as highlighted in [47, 48], the phase spectrum maintains global structure well, and the degradation of images is mainly manifested in the amplitude spectrum.
@@ III-B Overall Architecture
Let ${I}_{tar}$ and ${I}_{ref}\in\mathbb{R}^{H\times W}$ be a pair of MR images from the target modality and reference modality. Our objective is to reconstruct the high-quality target modality image $I_{tar}$ through a network that takes the under-sampled target modality image $\widetilde{I}_{tar}$ and the fully-sampled reference modality image ${I}_{ref}$ as inputs. Here, $\widetilde{I}_{tar}=\mathcal{F}^{-1}(\mathcal{M}\odot k_{tar})$, where $\mathcal{F}^{-1}$ represents the inverse Fourier transform (IFT), $\mathcal{M}$ denotes the binary under-sample mask, and $k_{tar}$ represents the fully-sampled k-space data of the target modality. The multi-modal reconstruction task can be formulated as the following optimization problem:
$\hat{I}_{tar}=\operatorname*{argmin}_{f(\theta)}\|I_{tar}-f(\mathcal{F}^{-1}(\mathcal{M}\odot k_{tar}),I_{ref};\theta)\|_{1}.$ (6)
To fully harness and fuse the information from different modalities, we design the MMR-Mamba network, depicted in Fig. 1. Initially, we adopt CNN blocks $\psi$ for low-level feature extraction and Mamba blocks $\phi$ for high-level feature extraction, producing modality-specific features $F_{tar}$ and $F_{ref}$ for the target modality and reference modality, respectively.
$F_{tar}=\phi_{tar}(\psi_{tar}(\widetilde{I}_{tar})),~{}F_{ref}=\phi_{ref}(\psi_{ref}({I}_{ref})).$ (7)
Subsequently, we design the Target modality-guided Cross Mamba (TCM) module in the spatial domain and the Selective Frequency Fusion (SFF) module in the frequency domain, generating domain-specific fused features $F_{spa}$ and $F_{fre}$:
$F_{spa}=TCM(F_{tar},F_{ref}),~{}F_{fre}=SFF(F_{tar},F_{ref}).$ (8)
Lastly, the multi-modal features in these two domains are further integrated through an Adaptive Spatial-Frequency Fusion (ASFF) module to generate the final fused feature, which is then fed into a CNN Decoder to reconstruct the target modality image:
@@
The model is trained by minimizing the $L_{1}$ loss between the reconstructed image and the ground-truth image:
$\mathcal{L}=\frac{1}{N}\sum_{n=1}^{N}\|\hat{I}_{tar}-I_{tar}\|_{1},$ (10)